# Entanglement phases and phase transitions in monitored free fermion systems of localization

### v2

@@ III.2 Phases and phase transitions
Figure 6: Finite-size scaling for the effective central charge with the BKT scenario, (a): $g(L)\gamma\bar{c}_{\rm eff}(\gamma,\Delta)$ versus $\log(L)-\nu/\sqrt{\gamma-\gamma_{c}}$, (b) : $g(L)\gamma\bar{c}_{\rm eff}(\gamma,V)$ versus $\log L-\nu/\sqrt{\gamma-\gamma_{c}}$ .(c): $g(L)\Delta\bar{c}_{\rm eff}(\gamma,\Delta)$ versus $\log(L)-\nu/\sqrt{\Delta-\Delta_{c}}$, (d) : $g(L)V\bar{c}_{\rm eff}(\gamma,V)$ versus $\log L-\nu/\sqrt{V-V_{c}}$. $g(L)=[1+1/(2\log L-\alpha)]^{-1}$ Detailed numerical data are provided in Table. 1.
@@
To quantitatively determine the location and nature of the entanglement phase transition, we perform a finite-size scaling and data collapse analysis of the effective entanglement quantity $\bar{c}_{\mathrm{eff}}$ within the Berezinskii-Kosterlitz-Thouless (BKT) framework. Given the pronounced finite-size corrections in free-fermion circuits, simple visual identification of crossing points is unreliable. We adopt a single-parameter scaling ansatz with a mild size correction factor $g(L)$, incorporating the essential singularity of the BKT correlation length, expressed as $g(L)\,x\,\bar{c}_{\mathrm{eff}}(x,y)=F[\log(L\,\xi(x,x_{c}(y)))]$, where $x\in\{\gamma,\Delta,V\}$ and $\xi(x,x_{c})\sim\exp[-\nu/\sqrt{|x-x_{c}|}]$. This yields natural scaling variables $X_{x}=\log L-\nu/\sqrt{|x-x_{c}|}$ and $Y_{x}=g(L)\,x\,\bar{c}_{\mathrm{eff}}$, allowing us to absorb leading finite-size drift and achieve data collapse onto a unified master curve across system sizes and parameter paths. Fig. 6 illustrates this systematic validation across multiple paths: (a) for fixed local potential $\Delta=(0.1,0.16,0.23,0.3)$, data from system sizes $L=80,96,112$ collapse well after rescaling, enabling precise determination of $\gamma_{c}(\Delta)$ and a shared $\nu$; (b) repeating the analysis for different potential strengths $V=\{0.3,0.8,1.3,1.8\}$, we trace the evolution of $\gamma_{c}(V)$; (c) scanning $\Delta$ at fixed measurement strengths $\gamma=\{0.16,0.23,0.28,0.30\}$, we extract $\Delta_{c}(\gamma)$, which agrees with the inverse function from (a); (d) scanning $V$ at fixed $\gamma=\{0.10,0.18,0.24,0.30\}$, we determine $V_{c}(\gamma)$, completing the critical boundary in the $(\gamma,V)$ plane. The critical points obtained from all four paths are consistently described using a unified scaling function $F$ and a common critical exponent $\nu$, indicating that the local potential does not merely shift an isolated critical point but cooperates with measurement strength to generate a continuous phase boundary in the $(\gamma,\Delta,V)$ parameter space. Mathematically, while measurement tends to suppress entanglement, the local potential restructures the state and alters information transport, and the competition between these effects determines the geometry and location of the phase boundary governed by the BKT singularity. The successful data collapses validate the applicability and robustness of the BKT scaling framework and provide a quantitative foundation for extracting the critical points $\gamma_{c}(\Delta)$, $\gamma_{c}(V)$, $\Delta_{c}(\gamma)$, $V_{c}(\gamma)$, and the correlation length exponent $\nu$.
@@
Furthermore, we present an understanding of the entanglement phase transition under the combined influence of localization potential and measurement: the phase boundary is jointly determined by both factors and depends on the type of localization induced by the potential, rather than the potential itself.
Our work highlights the distinctive behavior of MIPTs and relaxation dynamics in Stark-driven monitored systems, offers a new perspective on the characterization of the entanglement phase boundary arising from the combined effects of measurement and localization.
@@ Acknowledgements.
We thank Shuo Liu and Ze-Chuan Liu for the discussions. H.-Z. Li is supported by the CSC scholarship. J.-X. Zhong acknowledges the Shanghai University Distinguished Professor Research Start-up Project, the National Natural Science Foundation of China (Grant No.11874316 and No.12374046), the National Basic Research Program of China (Grant No.2015CB921103), and the Program for Changjiang Scholars and Innovative Research Team in University (Grant No.IRT13093).
@@ Appendix A Discussion on the Phase Boundary
After introducing measurements or local potentials, the entanglement growth of the system is modified. Denoting the corrections due to measurements and local potentials by $\Gamma(\gamma)$ and $\Gamma(P)$, respectively, we have:
$S(t)=\Gamma(\gamma)G_{0}t,S(t)=\Gamma(P)G_{0}t,$ (18)
Hence, once measurements and local potentials suppress the Entangling-phase growth rate down to that of the critical phase, the system reaches criticality:
$G_{0}\Gamma(\gamma)+G_{0}\Gamma(P)=C_{0}/(t+1),$ (19)
@@
$\max_{L,t}A(\gamma;L,t)\;\propto\;\exp\bigl{(}-\,a\,\gamma^{b}\bigr{)},$ (24)
@@
For the local potentials induced renormalization of the entanglement-growth rate $\Gamma(P)$, consider the single-particle tight-binding Hamiltonian as shown in Eq. (1), as an illustrative example, we choose the SP:
@@
Eq. (35) is the general solution we finally obtain, and in this paper we fix the parameters as $(a,b,c,d,e,C_{0})=(6,1.5,8,2,0.5,3)$.
@@ Appendix B Free fermions with Anderson disorder under measurement
We here complement the results for the SP and QPP by presenting the phase diagram of a one-dimensional chain of free fermions subject to Anderson-type disorder and continuous measurement. We consider spinless free fermions on a one-dimensional lattice with nearest-neighbor hopping $J$ in the presence of uncorrelated on-site disorder:
$H=-J\sum_{i=1}^{L-1}\left(c_{i}^{\dagger}c_{i+1}+c_{i+1}^{\dagger}c_{i}\right)+\sum_{i=1}^{L}\varepsilon_{i}\,n_{i},$ (36)
@@
where $c_{i}^{\dagger}$ ($c_{i}$) creates (annihilates) a fermion at site $i$, and $n_{i}=c_{i}^{\dagger}c_{i}$ is the local number operator. The on-site potential $\varepsilon_{i}$ is an independent random variable drawn from the uniform distribution $\left[-{W},{W}\right]$, where $W$ is the disorder strength. We set $J=1$ as the unit of energy and impose the same boundary conditions and filling as in the main text. In the Fig. 8, strong measurement or strong disorder yields an area-law phase with $\bar{c}_{\mathrm{eff}}\approx 0$, while weak measurement and weak disorder produce a logarithmic-entanglement phase with $\bar{c}_{\mathrm{eff}}>0$. The phase boundary predicted by Eq. (12) in the main text is determined by the competition between the measurement-induced disentangling rate and the effective propagation scale set by localization, and it agrees very well with the numerically obtained transition line. Using the same parameters as for the quasiperiodic potential, we achieve an accurate fit to the phase boundary in the disordered phase diagram. This further supports our conclusion that the phase boundary depends only on the localization characteristics of the wave function; that is, as long as the lattice potential leads to the same degree of localization, the phase boundaries will be similar, largely independent of the specific form of the potential used in the model.

### v4

@@ III.2 Phases and phase transitions
Figure 6: Finite-size scaling for the effective central charge with the BKT scenario, (a): $g(L)\gamma\bar{c}_{\rm eff}(\gamma,\Delta)$ versus $\log(L)-\nu/\sqrt{\gamma-\gamma_{c}}$, (b) : $g(L)\gamma\bar{c}_{\rm eff}(\gamma,V)$ versus $\log L-\nu/\sqrt{\gamma-\gamma_{c}}$ .(c): $g(L)\Delta\bar{c}_{\rm eff}(\gamma,\Delta)$ versus $\log(L)-\nu/\sqrt{\Delta-\Delta_{c}}$, (d) : $g(L)V\bar{c}_{\rm eff}(\gamma,V)$ versus $\log L-\nu/\sqrt{V-V_{c}}$. $g(L)=[1+1/(2\log L-\alpha)]^{-1}$. Detailed numerical data are provided in Table. 1.
To quantitatively determine the location and nature of the entanglement phase transition, we perform a finite-size scaling and data collapse analysis of the effective entanglement quantity $\bar{c}_{\mathrm{eff}}$ within the Berezinskii-Kosterlitz-Thouless (BKT) framework. Given the pronounced finite-size corrections in free-fermion circuits, simple visual identification of crossing points is unreliable. We adopt a single-parameter scaling ansatz with a mild size correction factor $g(L)$, incorporating the essential singularity of the BKT correlation length, expressed as [6, 44]
$g(L)\,x\,\bar{c}_{\mathrm{eff}}(x,y)=F[\log(L\,\xi(x,x_{c}(y)))],$ (16)
Figure 7: The finite-size scaling collapse of the entanglement entropy is performed under the assumption of BKT scaling of the correlation length. Subfigures (a) and (b) correspond respectively to the critical points $(\gamma_{c},\Delta_{c})=(0.16,0.3),(0.23,0.23),(0.28,0.16),(0.30,0.10)$, where panel (a) is obtained from the collapse of $\gamma$ data and panel (b) from the collapse of $\Delta$ data. Subfigures (c) and (d) correspond respectively to the critical points $(\gamma_{c},V_{c})=(0.10,1.8),(0.18,1.3),(0.24,1.8),(0.30,0.3)$, where panel (c) is obtained from the collapse of $\gamma$ data and panel (d) from the collapse of $V$ data. The analyses are performed using system sizes $L=12,24,48,\ldots,112$.
Table 1: Critical exponents and critical points.
$\Delta$ $\gamma_{c}$ $\nu$ $\alpha$
0.10 $0.30\pm 0.035$ $3.8\pm 0.2$ $5.9\pm 0.3$
0.16 $0.28\pm 0.035$ $4.1\pm 0.2$ $6.2\pm 0.3$
0.23 $0.23\pm 0.045$ $5.7\pm 0.3$ $6.5\pm 0.45$
0.30 $0.16\pm 0.040$ $6.5\pm 0.3$ $7.3\pm 0.5$
$V$ $\gamma_{c}$ $\nu$ $\alpha$
0.30 $0.30\pm 0.04$ $2.7\pm 0.3$ $4.1\pm 0.35$
0.80 $0.24\pm 0.04$ $2.9\pm 0.35$ $4.3\pm 0.35$
1.30 $0.18\pm 0.04$ $3.3\pm 0.35$ $4.5\pm 0.4$
1.80 $0.10\pm 0.045$ $3.8\pm 0.4$ $4.9\pm 0.4$
$\gamma(\Delta)$ $\Delta_{c}$ $\nu$ $\alpha$
0.16 $0.30\pm 0.015$ $6.5\pm 0.1$ $8.9\pm 0.1$
0.23 $0.23\pm 0.03$ $4.9\pm 0.2$ $7.4\pm 0.3$
0.28 $0.16\pm 0.035$ $4.1\pm 0.3$ $6.5\pm 0.4$
0.30 $0.10\pm 0.035$ $3.3\pm 0.3$ $4.9\pm 0.4$
$\gamma(V)$ $V_{c}$ $\nu$ $\alpha$
0.10 $1.8\pm 0.35$ $3.6\pm 0.35$ $5.3\pm 0.45$
0.18 $1.3\pm 0.3$ $3.2\pm 0.3$ $4.8\pm 0.45$
0.24 $0.8\pm 0.15$ $2.9\pm 0.2$ $4.1\pm 0.3$
0.30 $0.1\pm 0.28$ $2.6\pm 0.3$ $3.2\pm 0.4$
@@
where $x\in\{\gamma,\Delta,V\}$ and correlation length $\xi(x,x_{c})\sim\exp[-\nu/\sqrt{|x-x_{c}|}]$. This yields natural scaling variables $X_{x}=\log L-\nu/\sqrt{|x-x_{c}|}$ and $Y_{x}=g(L)\,x\,\bar{c}_{\mathrm{eff}}$, allowing us to absorb leading finite-size drift and achieve data collapse onto a unified master curve across system sizes and parameter paths. Fig. 6 illustrates this systematic validation across multiple paths: (a) for fixed localized potential $\Delta=(0.1,0.16,0.23,0.3)$, data from system sizes $L=80,96,112$ collapse well after rescaling, enabling precise determination of $\gamma_{c}(\Delta)$ and a shared $\nu$; (b) repeating the analysis for different potential strengths $V=\{0.3,0.8,1.3,1.8\}$, we trace the evolution of $\gamma_{c}(V)$; (c) scanning $\Delta$ at fixed measurement strengths $\gamma=\{0.16,0.23,0.28,0.30\}$, we extract $\Delta_{c}(\gamma)$, which agrees with the inverse function from (a); (d) scanning $V$ at fixed $\gamma=\{0.10,0.18,0.24,0.30\}$, we determine $V_{c}(\gamma)$, completing the critical boundary in the $(\gamma,V)$ plane. These appropriate values of $\gamma_{c}$ or $\Delta_{c},V_{c}$ lead to good scaling collapses, which are consistent with the scaling formula Eq. 16. To more precisely determine the phase transition points and to examine whether the transition belongs to the BKT universality class, we perform finite-size scaling analyses for the four critical points using the following scaling relations [44]:
$\bar{S}_{\gamma}-\bar{S}_{\gamma_{c}}$ $=F\!\left[(\gamma-\gamma_{c})(\log L)^{2}\right],$ (17)
$\bar{S}_{\Delta}-\bar{S}_{\Delta_{c}}$ $=F\!\left[(\Delta-\Delta_{c})(\log L)^{2}\right],$
$\bar{S}_{V}-\bar{S}_{V_{c}}$ $=F\!\left[(V-V_{c})(\log L)^{2}\right],$
where $F$ denotes the scaling function. This scaling formula is based on the assumption that the MIPT belongs to a universality class similar to the BKT one [6], where the correlation length $\xi$ diverges exponentially around the transition point; specifically,
$\log\xi$ $\sim\frac{1}{\sqrt{\gamma-\gamma_{c}}}\quad$ $\text{as }\gamma\to\gamma_{c}+0,$ (18)
$\log\xi$ $\sim\frac{1}{\sqrt{\Delta-\Delta_{c}}}\quad$ $\text{as }\Delta\to\Delta_{c}+0,$
$\log\xi$ $\sim\frac{1}{\sqrt{V-V_{c}}}\quad$ $\text{as }V\to V_{c}+0.$
As shown in Fig. 7, these transition points yield good scaling collapses, consistent with Eq. 17. The successful scaling collapses provide strong evidence that the phase transition, jointly driven by the measurement strength $\gamma$ and the Stark potential strength $\Delta$, falls into the BKT universality class. To further determine the error of the phase transition point , we minimize the cost function [102], where the error bar of $\gamma_{c}$ or $\Delta_{c},V_{c}$ are estimated from the range within which the cost function does not exceed twice their minimum value. The obtained results are presented as circles with error bars in Fig. 2 and are summarized in Table. 1. The critical points obtained from all four paths are consistently described using a unified scaling function $F$ and a common critical exponent $\nu$, indicating that the localized potential does not merely shift an isolated critical point but cooperates with measurement strength to generate a continuous phase boundary in the $(\gamma,\Delta,V)$ parameter space. Mathematically, while measurement tends to suppress entanglement, the localized potential restructures the state and alters information transport, and the competition between these effects determines the geometry and location of the phase boundary governed by the BKT singularity. The successful data collapses validate the applicability and robustness of the BKT scaling framework and provide a quantitative foundation for extracting the critical points $\gamma_{c}(\Delta)$, $\gamma_{c}(V)$, $\Delta_{c}(\gamma)$, $V_{c}(\gamma)$, and the correlation length exponent $\nu$.
@@
Furthermore, we construct a phase diagram under different localization mechanisms, which diverges at a boundary characterized by entanglement propagation. We find that the monitored Stark-localized system decays rapidly under weak measurement strength, whereas the Anderson-localized system exhibits a slower decay.
Our work highlights the distinctive behavior of MIPTs and relaxation dynamics in Stark-driven monitored systems, and provides a perspective on the characterization of the entanglement phase boundary arising from the combined effects of measurement and localization.
@@ Acknowledgements.
We thank Shuo Liu and Ze-Chuan Liu for the discussions. J.-X. Zhong acknowledges the National Natural Science Foundation of China (Grant No.12374046) and the Shanghai Science and Technology Action Plan (Grant No.24LZ1400800). H.-Z. Li is supported by the CSC scholarship.
@@ Appendix A Discussion on the Phase Boundary
After introducing measurements or localized potentials, the entanglement growth of the system is modified. Denoting the corrections due to measurements and localized potentials by $\Gamma(\gamma)$ and $\Gamma(P)$, respectively, we have:
$S(t)=\Gamma(\gamma)G_{0}t,S(t)=\Gamma(P)G_{0}t.$ (21)
Hence, once measurements and localized potentials suppress the Entangling-phase growth rate down to that of the critical phase, the system reaches criticality:
$G_{0}\Gamma(\gamma)+G_{0}\Gamma(P)=C_{0}/(t+1).$ (22)
@@
$\max_{L,t}A(\gamma;L,t)\;\propto\;\exp\bigl(-\,a\,\gamma^{b}\bigr).$ (27)
@@
For the localized potentials induced renormalization of the entanglement-growth rate $\Gamma(P)$, consider the single-particle tight-binding Hamiltonian as shown in Eq. (1), as an illustrative example, we choose the SP:
@@
Eq. (38) is the general solution we finally obtain, and in this paper we fix the parameters as $(a,b,c,d,e,C_{0})=(6,1.5,8,2,0.5,3)$, these parameters are not fundamental physical constants derived analytically, but rather representative values chosen through empirical fitting. The analytical analysis only determines the envelope function form describing the influence of the measurement strength and the localized potential strength on the entanglement growth rate, while the explicit numerical values of the parameters cannot be obtained in closed form. The selected set therefore serves as an effective choice that provides a reasonable description of the numerical phase boundaries and allows us to generate the plots.
@@ Appendix B Free fermions with Anderson disorder under measurement
We here complement the results for the SP and QPP by presenting the phase diagram of a one-dimensional chain of free fermions subject to Anderson-type disorder and continuous measurement. We consider spinless free fermions on a one-dimensional lattice with nearest-neighbor hopping $J$ in the presence of uncorrelated on-site disorder as shown in Eq. (1), We set $J=1$ as the unit of energy and impose the same boundary conditions and filling as in the main text. In the Fig. 9, strong measurement or strong disorder yields an area-law phase with $\bar{c}_{\mathrm{eff}}\approx 0$, while weak measurement and weak disorder produce a logarithmic-entanglement phase with $\bar{c}_{\mathrm{eff}}>0$. The phase boundary predicted by Eq. (12) in the main text is determined by the competition between the measurement-induced disentangling rate and the effective propagation scale set by localization, and it agrees very well with the numerically obtained transition line. Using the same parameters as for the quasiperiodic potential, we achieve an accurate fit to the phase boundary in the disordered phase diagram. This further supports our conclusion that the phase boundary depends only on the localization characteristics of the wave function; that is, as long as the lattice potential leads to the same degree of localization, the phase boundaries will be similar, largely independent of the specific form of the potential used in the model.
Appendix C Explanation of the measurement process
We are monitoring the local particle number operator $n_{j}=c_{j}^{\dagger}c_{j}$ through continuous weak measurement as shown in Fig. 10 .
Figure 10: When a projective measurement is performed on an arbitrary quantum state, the system collapses into an eigenstate of the measurement operator. Consequently, repeated measurements of the same observable will necessarily yield identical outcomes. In contrast, a weak measurement is implemented by weakly coupling the system to an ancillary qubit $\phi$, which is then read out. This procedure extracts only partial information about the system without fully projecting it onto an eigenstate. Therefore, a subsequent measurement of the same observable may yield a different outcome and leave the system in a different post-measurement state.
Specifically, different from projective measurement, the core idea of weak measurement is to employ a weak coupling between the system and the measurement apparatus, such that each individual measurement only slightly perturbs the system state while still extracting partial information. To formalize this process, we introduce a continuous measurement outcome $x$ and construct a series of Kraus operators:
$M(x)=\frac{1}{\left(2\pi\sigma^{2}\right)^{1/4}}e^{-\frac{(x-\omega\hat{O}dt)^{2}}{4\sigma^{2}}},$ (39)
where, $\sigma$ describes the width of the probe initial wave function. These operators satisfy the continuous completeness relation $\int dxM^{\dagger}(x)M(x)=I$. For a single measurement on an initial state $|\psi\rangle$, the probability density of obtaining the result $x$ is given by $p(x)=\langle\psi|M^{\dagger}(x)M(x)|\psi\rangle$, and the normalized post-measurement state is updated to
$|\psi(x)\rangle=\frac{M(x)|\psi\rangle}{\sqrt{p(x)}}.$ (40)
In Kraus representation, to track the evolution of a monitored quantum system at the pure state level and capture the randomness introduced by measurements, one often employs the quantum trajectory method, that is, the Stochastic schr¨odinger equation. The evolution of the system over a very short time interval $dt$ is represented by a set of Kraus operators $\left\{M_{\alpha}\right\}$, so that the density matrix evolves as:
$\rho(t+dt)=\sum_{\alpha}M_{\alpha}(dt)\rho(t)M_{\alpha}^{\dagger}(dt).$ (41)
These operators can be divided into two categories:
$M_{0}=I-iH_{\mathrm{eff}}dt,\quad M_{k}=\sqrt{\gamma dt}L_{k}.$ (42)
Here, $H_{\text{eff }}=H-\frac{i\gamma}{2}\sum_{k\geq 1}L_{k}^{\dagger}L_{k}$ is the non-Hermitian effective Hamiltonian and $L_{k}$ is the jump operator. $M_{0}$ and $M_{k}$ denote the no-jump operator (sometimes referred to as the non-Hermitian evolution operator) and the jump operator, respectively. In the discrete picture, during each small time step $dt$, the system either evolves without a jump or experiences a single jump. The probability for a jump is determined by a corresponding expression
$p_{k\geq 1}=\langle\psi(t)|M_{k}^{\dagger}M_{k}|\psi(t)\rangle\approx\gamma dt\langle\psi(t)|L_{k}^{\dagger}L_{k}|\psi(t)\rangle.$ (43)
If a jump occurs, the state vector is updated according to the jump operator, i.e.,
$|\psi(t+dt)\rangle=\frac{M_{k}|\psi(t)\rangle}{\sqrt{p_{k\geq 1}}},$ (44)
and if no jump occurs, the state vector evolves under the no-jump operator:
$|\psi(t+dt)\rangle=\frac{M_{0}|\psi(t)\rangle}{\sqrt{p_{0}}},$ (45)
where $p_{0}=1-\sum_{k\geq 1}p_{k}$ is the probability for no jump. This framework effectively captures the stochastic nature of the evolution in a monitored quantum system.
In the continuous limit $dt\rightarrow 0$, the evolution process can be described by an Itô stochastic differential equation. One of the most common formulations is known as the quantum jump unraveling. In this approach, the state evolves continuously under a non-Hermitian effective Hamiltonian until a quantum jump occurs. The probability of a jump in an infinitesimal time interval is determined by the jump operator, and when a jump occurs, the state is updated according to that operator. Between jumps, the state follows a deterministic (but non-unitary) evolution. Formally, the quantum jump stochastic Schrödinger equation can be written as
$d|\psi(t)\rangle$ $=-iH_{\mathrm{eff}}|\psi(t)\rangle dt$
$\quad+\sum_{k\geq 1}\left(\frac{L_{k}|\psi(t)\rangle}{\|L_{k}|\psi(t)\rangle\|}-|\psi(t)\rangle\right)dN_{k}(t),$ (46)
where $dN_{k}(t)$ is the set of Poisson stochastic increment, this is the Eq. (II) and the Born rule is indeed encoded in the norm of the jump operators $\|L_{k}|\psi(t)\rangle\|$. The Poisson increments $dN_{k}(t)$ stochastically determine whether a jump occurs in the interval $dt$ (in the simulations, we set $dt=0.05$. ). The probability for a jump is given by
$p_{k}=\langle\psi(t)|L_{k}^{\dagger}L_{k}|\psi(t)\rangle dt=\|L_{k}|\psi(t)\rangle\|^{2}dt,$ (47)
which is precisely the Born probability. The no-jump probability is $p_{0}=1-\sum_{k}p_{k}$. Thus, the randomness of each trajectory is fully governed by the Born rule, not by any artificial filtering. Physical observables are then obtained by averaging over the full ensemble of trajectories, rather than by keeping only a specific subset.